# Differentially private and decentralized randomized power method

### v1

@@ 6 Decentralized version
In this section, we explore the scenario where the matrix $\bm{A}$ is distributed across multiple clients. Specifically, each client holds a matrix $\bm{A}^{(1)},\cdots,\bm{A}^{(s)}\in\mathbb{R}^{n\times n}$ such that $\sum_{i=1}^{s}\bm{A}^{(s)}=\bm{A}$. The goal is for the clients to collaboratively compute an orthonormal basis for the range of $\bm{A}$, similar to the centralized randomized power method, but without revealing their individual private matrices $\bm{A}^{(s)}$ to the server or other clients.
The randomized power method involves linear operations, which makes it well-suited for parallelization. An implementation of a private and distributed power method, using public channel communication between the clients and the server, was proposed in (Balcan et al., 2016). However, this approach relies on local Differential Privacy, since the data exchanged between the clients and the server can be observed by everyone.
To improve privacy while keeping the benefits of distributed computation, we propose using Secure Aggregation, a lightweight Multi-Party Computation (MPC) protocol. Secure Aggregation eliminates the need for local DP by allowing us to use distributed DP, which offers privacy guarantees similar to central DP. This approach has been extensively studied in the literature (Chen et al., 2021; Kairouz et al., 2021; Goryczka et al., 2013; Wei et al., 2024; Ghazi et al., 2019).
In distributed DP, each client adds noise to its own contributions before using the Secure Aggregation protocol. The noise is calibrated such that the sum of the noise across clients has a variance comparable to that used in Central DP, thus achieving similar privacy guarantees without requiring a trusted curator to aggregate the data.
We assume the “honest-but-curious” threat model, meaning that while the clients may attempt to learn additional information, they follow the protocol honestly. We also assume no corrupted or dropout users. For simplicity, this paper focuses on a basic version of distributed DP, neglecting effects of data quantization or errors introduced by modular arithmetic. We refer to (Kairouz et al., 2021) for an implementation of Distributed DP that takes these effects into account.
Finally, it is also possible to use a fully decentralized and differentially private protocol to perform averaging (Sabater et al., 2022). It is therefore straightforward to transform our protocol to avoid relying on a central server for Secure Aggregation, while keeping similar utility and communication costs.
@@ 6.1 Decentralized and private power method using distributed DP:
We give a federated version of the PPM in Algorithm 2. In this version, we reduce the variance of the noise by a multiplicative factor in O($s\cdot p\log({d})\cdot\|\bm{U}\|_{\infty}^{2}$) compared to (Balcan et al., 2016) due to the fact that we emulate Central DP with Secure Aggregation (thus using a secure channel) rather than using a public channel that only allows for local DP, and because we propose a generally tighter sensitivity bound. We show the equivalence between Algorithm 2 and Algorithm 1 in Theorem 6.2, and provide the proof in Appendix A.5. We first define the secure aggregation algorithm.
@@
Theorem 6.2 (Privacy and convergence of Algorithm 2).
Algorithm 2 is as privacy-preserving and has the same utility as its centralized version (Algorithm 1).
@@ 6.2 Communication and computation cost analysis:
7 Conclusion
@@
We presented a differentially private and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems. We introduced new convergence bounds, improving the accuracy of the method while ensuring privacy guarantees. By employing Secure Aggregation in a decentralized setting, we were able to reduce the noise introduced for privacy, leading to higher utility without compromising on privacy. The decentralized variant of the algorithm maintains the efficiency and privacy of the centralized version but adapts it for distributed environments. Our theoretical and empirical results demonstrate the utility and privacy advantages of our approach, suggesting that it offers a strong solution for real-world decentralized applications.
8 Limitations
@@
The proposed method is effective in ensuring differential privacy and decentralization but still has limitations. The privacy guarantees rely on the honest-but-curious threat model and assume no collusion or malicious behavior from participants, which may not always be realistic in certain environments. In the future, it would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially have a faster convergence.
@@ Appendix A Proofs of Results
$\begin{split}\Delta_{l}&=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}\\ &\leq\sqrt{\sum_{i}(\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}(\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\triangleq\Delta_{l}^{b}.\\ \end{split}$ (24)
Although the proposed update model (12) is more general than when $\bm{A}^{\prime}$ is defined using (11), the proposed bound $\Delta_{l}^{b}$ is also generally tighter than the bound proposed in (Hardt & Price, 2014; Balcan et al., 2016).
A.2 Proof of Theorem 3.2
@@
$\leq\sum_{i=1}^{L}\frac{(\Delta_{l}^{b})^{2}}{2\sigma_{i}^{2}}$ (26)
@@
By Proposition 1.3 of (Bun & Steinke, 2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$. Let $\epsilon^{\prime}\triangleq\rho+2\sqrt{\rho\log(1/\delta)}$.
@@
Which leads us to the following Theorem by injecting $\sigma_{l}=\Delta_{l}^{b}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$.
A.4 Derivation of Approximation 5.1
We provide here the derivation for Approximation 5.1:
@@
Let $\bm{M}\in\mathbb{R}^{n\times k}$ with i.i.d $\mathcal{N}(0,1)$-distributed entries. Let $\bm{M}=\bm{Q}\bm{R}$ be its QR factorization (by definition $\bm{i}sQ$ orthogonal and $\bm{R}$ upper triangular). By the Barlett decomposition theorem (Muirhead, 2009), we know that $\bm{Q}$ is a random matrix distributed uniformly in the Stiefel manifold $\mathbb{V}_{k,n}$. Then by Theorem 2.2.1 of (Chikuse, 2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries. We approximate $\bm{Z}^{\top}\bm{Z}$ as a diagonal matrix and remark that its diagonal elements are distributed as random chi-squared variables with $n$ degrees of liberty. For a matrix $\bm{A}$, we note as $\bm{A}^{\circ n}$ the elementwise Hadamard exponentiation. Then we have:
@@
We can approximate the expectancy of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 of (Massart, 1896). We therefore get approximations of $\mathbb{E}[\max_{ij}k\bm{Q}_{ij}^{2}]$ and $\mathbb{E}[\max_{i}\|\bm{Q}_{i:}\|^{2}]$. We assume that the variances of $\Delta_{l}^{a}$ and $\Delta_{l}^{b}$ (as $n$ and $k$ are large) are small and therefore:
$\mathbb{E}[\Delta_{l}^{a}]^{2}\approx\mathbb{E}[(\Delta_{l}^{a})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}$ (46)
$\mathbb{E}[\Delta_{l}^{b}]^{2}\approx\mathbb{E}[(\Delta_{l}^{b})^{2}]=\sqrt{\frac{k\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(n)}+\frac{k\mu}{n}.$ (47)
@@
First bound: We know by Section 2.4 of (Woodruff et al., 2014) that any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$ has squared norm at most 1. Therefore $\max_{i}\|\bm{X}_{i:}\|_{2}\leq 1$ because $\bm{X}$ has orthonormal columns by construction.
Second bound: Without loss of generality and to simplify notation, we note $\bm{X}$ to denote a matrix $\bm{X}_{l}$ computed during the PPM. As proposed in (Hardt & Price, 2014), we can write any column $x_{c}$ of $\bm{X}$ ($\bm{X}_{:c}$) as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$ where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar. Therefore we have:

### v2

@@ 5 Decentralized version
In this section, we consider a decentralized setting where the matrix $\bm{A}$ is distributed across multiple clients. Specifically, each client $i$ holds a private matrix $\bm{A}^{(i)}\in\mathbb{R}^{n\times n}$, such that the global matrix is the sum of these local matrices: $\bm{A}=\sum_{i=1}^{s}\bm{A}^{(i)}.$
The goal is for the clients to collaboratively compute an orthonormal basis for the range of $\bm{A}$, similar to the centralized randomized power method, but without revealing their individual private matrices $\bm{A}^{(i)}$ to the server or to other clients.
The randomized power method involves linear operations, making it well-suited for parallelization and distributed computation. Balcan et al. (2016) proposed a private and federated power method using communication over public channels between clients and a server. However, this approach relies on Differential Privacy, since the data exchanged can be observed by everyone, which requires high levels of noise to ensure privacy.
To enhance privacy while retaining the benefits of distributed computation, we use Secure Aggregation, a lightweight Multi-Party Computation protocol. Secure Aggregation allows clients to collaboratively compute sums without revealing individual data, enabling the use of distributed DP. This approach offers privacy guarantees similar to central DP and eliminates the need for a trusted curator. Distributed DP has been extensively studied in the literature (Chen et al., 2021; Kairouz et al., 2021; Goryczka et al., 2013; Wei et al., 2024; Ghazi et al., 2019).
In distributed DP, each client adds carefully calibrated noise to their local contributions before participating in the Secure Aggregation protocol. The noise is designed such that the sum across clients has a variance comparable to that used in central DP, thereby achieving similar Differential Privacy guarantees without requiring a trusted aggregator, and allowing clients to keep the data locally.
We operate under the “honest-but-curious“ threat model, where clients follow the protocol correctly but may attempt to learn additional information from received data. We also assume there are no corrupted or dropout users during the computation. For simplicity, this paper focuses on a basic implementation of distributed DP, neglecting the effects of data quantization or errors introduced by modular arithmetic. We refer to Kairouz et al. (2021) for an implementation of Distributed DP that takes these effects into account.
Finally, it is possible to use a fully decentralized and differentially private protocol to perform averaging without relying on a central server for Secure Aggregation, as demonstrated by Sabater et al. (2022). Adopting such approaches can further enhance decentralization while maintaining similar utility and communication costs.
@@ 5.1 Decentralized and private power method using distributed DP:
We introduce a federated version of the Privacy-Preserving Power Method in Algorithm 2. This version reduces the noise variance by a factor of $O(sp\log d\,\|\bm{U}\|_{\infty}^{2})$ compared to the method in Balcan et al. (2016). The improvement comes from the fact that we emulate centralized Differential Privacy (DP) over secure channels, avoiding the higher noise required in local DP settings with public channels, and that we use our generally tighter sensitivity bound.
We demonstrate the equivalence between the decentralized Algorithm 2 and the centralized Algorithm 1 in Theorem 5.2 (proof provided in Appendix A.5). We also give a simplified definition of the Secure Aggregation algorithm.
@@
Theorem 5.2 (Privacy and utility of Algorithm 2).
The decentralized Privacy-Preserving Power Method (Algorithm 2) provides the same privacy guarantees and achieves equivalent utility (in terms of convergence) as its centralized version (Algorithm 1).
@@ 5.2 Communication and computation cost analysis:
6 Empirical comparison of the proposed bounds
We introduced both runtime-dependent and runtime-independent convergence bounds for our algorithm. The runtime-dependent bound depends on the matrix $\bm{A}$ and the iteratively computed matrices $\bm{X}_{l}$, whereas the runtime-independent bound (Theorem 4.1) only depends on $\bm{A}$. To illustrate the practical impact of our changes, we focus on an application in recommender systems and compare how our algorithm perform in this context. We will then use a statistical approximation to see how the runtime-dependent bounds compare at the first step of the algorithm, regardless of the application.
6.1 Application to recommender systems
State-of-the-art recommender systems like GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) utilize singular value decomposition (SVD) or eigenvalue decomposition as part of their algorithms. Specifically, they represent the dataset (comprised of user-item interactions) using a normalized adjacency matrix $\tilde{\bm{R}}$, from which they compute a normalized item-item matrix $\tilde{\bm{P}}$. To get rid of noise in $\tilde{\bm{R}}$, they compute an ideal low-pass filter based on the top-$p$ eigenvectors of $\tilde{\bm{P}}$ and apply it to $\tilde{\bm{R}}$. Detailed definitions are provided in Appendix A.7.
Table 2 presents key statistics of popular recommendation datasets used in the literature. Depending on the use case, these datasets may be held by a central curator, or distributed among the users of a recommender system. Accordingly, either the centralized PPM (Algorithm 1) or the decentralized PPM (Algorithm 2) can be applied to compute the leading eigenvectors. To demonstrate the practicality of our proposed methods, we focus on a decentralized setting where each user has access to part of the dataset and collaborate with other users to compute the desired eigenvectors.
Table 2: Statistics of datasets
Dataset Users Items Interactions $\mu_{0}$ $\mu_{1}$ $\Delta_{l}^{prior}$ (existing) ${\hat{\Delta}_{l}}$ (ours)
Amazon-book 52,643 91,599 2,984,108 0.33 52.95 1.12 ×$\times\sqrt{p}$ 1
MovieLens 71,567 10,677 7,972,582 0.39 32.84 1.19 ×$\times\sqrt{p}$ 1
EachMovie 74,425 1,649 2,216,887 0.49 21.89 1.33 ×$\times\sqrt{p}$ 1
Jester 54,906 151 1,450,010 0.73 7.78 1.64 ×$\times\sqrt{p}$ 1
6.1.1 Empirical comparison of the runtime-independent bounds
As discussed in Section 4.1, our proposed runtime-independent bound is tighter than those proposed in Hardt & Price (2014); Balcan et al. (2016) when $\min(1,\mu_{1}(\bm{A}))\leq\mu_{0}(\bm{A})\sqrt{p\cdot\log({d})}$, where $\bm{U}$ denotes the eigenvectors of $\bm{A}$. Table 2 shows that this condition holds for popular recommender system datasets. As we can deduce from Table 2, the proposed bound theoretically allows us to converge to solutions with much smaller $\eta$ values, especially when the desired number of factors $p$ is large.
6.1.2 Practical utility of the proposed algorithms
We saw that the proposed runtime-independent bounds were practically tighter than the previous one for the task of interest. We demonstrate that our decentralized PPM can privately compute the top-$p$ eigenvectors $\bm{U}_{p}$ of the user-normalized item-item matrix $\tilde{\bm{P}}$ (proof in Appendix A.1):
Lemma 6.1.
Algorithm 2 with $\Delta_{l}=\sqrt{2}\max_{i}\|\bm{X}_{i:}^{l}\|_{2}$ can compute $\bm{U}_{p}$, the top-$p$ eigenvectors of $\tilde{\bm{P}}$ in a decentralized setting under $(\epsilon,\delta)$-Differential Privacy.
Figure 1: Impact of the Differential Privacy parameter $\epsilon$ on the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to multiple PPM versions, computed with $p=32$ and $L=3$ on the EachMovie dataset.
Both GF-CF (Shen et al., 2021) and BSPM (Choi et al., 2023) use $\bm{U}_{p}$ to compute the ideal low pass filter and apply it to the interaction matrix $\bm{R}$, giving a filtered interaction matrix $\bm{R}_{p}$. To illustrate the practicality of our proposed method, we compute approximations of $\bm{U}_{p}$ (denoted by $\tilde{\bm{U}}_{p}$) using either our proposed decentralized PPM, or the PPM versions proposed by  Balcan et al. (2016). We then use $\tilde{\bm{U}}_{p}$ to compute an approximation of $\bm{R}_{p}$ (noted $\tilde{\bm{R}_{p}}$), and compute the relative approximation error $\frac{\|\tilde{\bm{R}_{p}}-\bm{R}_{p}\|_{F}}{\|\bm{R}_{p}\|_{F}}$ associated to each version.
Figure 1 illustrates the impact of the privacy parameter $\epsilon$ on the approximation error for the EachMovie dataset, with $p=32$ and $L=3$. Our decentralized PPM achieves a relatively low approximation error (approximately 1/10 of the random guessing error) for practical values of $\epsilon$ (5-10). In contrast, the methods from Balcan et al. (2016) require $\epsilon$ values of at least $10^{3}$ to achieve comparable errors, which does not provide meaningful privacy protection. Additional comparisons and details are provided in Appendix A.7.1.
6.2 Matrix-agnostic comparison of the runtime-dependent bounds
We analyze the tightness of our proposed runtime-dependent bound compared to the previous bound in a matrix-agnostic manner, focusing on the first iteration of the algorithm.
At iteration $l=0$, the leading singular vectors are initialized as $\bm{X}^{0}=\text{Q}(\bm{\Omega})$, where $\bm{\Omega}\sim\mathcal{N}(0,1)^{n\times k}$ and $\text{Q}(\bm{\Omega})$ denotes the orthonormal matrix obtained from the QR decomposition of $\bm{\Omega}$. Since $\bm{X}^{0}$ is independent of the matrix $\bm{A}$, it allows us to assess the relative tightness of the proposed bound $\hat{\Delta}_{0}$ compared to the previous bound $\Delta_{0}^{prior}$ at the first step.
Let $r(k,n)$ be the ratio of the expected values of the two bounds, as a function of the desired number of eigenvectors $k$ and the matrix dimension $n$:
$r(k,n)=\frac{\mathbb{E}[\Delta_{0}^{prior}]}{\mathbb{E}[\hat{\Delta}_{0}]}.$ (18)
6.2.1 Theoretical approximation of $r(k,n)$:
We derive an asymptotic approximation of $r(k,n)$ using Approximation 6.2 (detailed in Appendix A.4).
Approximation 6.2.
[…]
$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(kn)\frac{k^{2}}{n^{2}}}+\frac{k\mu}{n}\right)$ (19)
$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\left(\sqrt{2\sigma^{2}\cdot\log(n)\frac{k}{n^{2}}}+\frac{k\mu}{n}\right).$ (20)
6.2.2 Empirical approximation of $r(k,n)$:
To compare the tightness of our proposed bound $\hat{\Delta}_{0}$ to the previous bound $\Delta_{0}^{prior}$, we perform both theoretical and empirical analyses at the first iteration, where the noise scaling depends only on the random initialization $\bm{X}^{0}$. Since $\bm{X}^{0}$ is independent of $\bm{A}$, we can estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$ by sampling random Gaussian matrices $\bm{X}^{0}$.
Experiment: $r(k,n)$ depends on the number of factors $k$ and on $n$, where ($n\times n$) is the size of $\bm{A}$. We therefore seek to compute it for multiple values of $n$ ($n\in\{8000,12000\}$) and a range of values for $k$ (between 64 and 4000, with steps of 64), for each value of $n$. We use Theorem 6.2 to approximate it theoretically. To measure it empirically, we compute $\Delta_{0}^{prior}$ and $\hat{\Delta}_{0}$ based on $\bm{X}_{0}$ in the first step of the algorithm (it is independent of the matrix of interest $\bm{A}$), for $t$ runs of the algorithm. We can then use the $t$ measures to estimate $\mathbb{E}[\Delta_{0}^{prior}]$ and $\mathbb{E}[\hat{\Delta}_{0}]$. We claim that $t=5$ is enough to see a general tendancy, as we have estimates for many ($k,n$) couples. We can also compare the empirical estimates to the proposed asymptotic approximations specified in Approximation 6.2.
Figure 2 and Figure 3 (in Appendix) presents the comparison between empirical (blue dots) and theoretical (green line) estimates of $r(k,n)$ for $n=8000$ and $n=12000$. The empirical ratios are calculated from the averages over the $t=5$ runs for each value of $k$. The results indicate that both empirically and theoretically the proposed noise scaling $\hat{\Delta}_{l}$ is much tighter than $\Delta_{l}^{prior}$ at the first step of the algorithm. Our theoretical approximation is conservative and underestimates how much tighter the bound is initially, compared to what we observe in practice. The proposed bound is tighter by a multiplicative factor on the first step and can therefore drastically reduce the impact of the noise introduced by DP on the convergence of the Privacy Preserving power method, as the algorithm is usually run for very few steps ($L$ is usually in the range of $1$-$5$).
7 Conclusion and limitations
@@
We presented a differentially private and decentralized randomized power method that addresses privacy concerns in large-scale spectral analysis and recommendation systems. We introduced new convergence bounds, improving the accuracy of the method while ensuring privacy guarantees. By employing Secure Aggregation in a decentralized setting, we were able to reduce the noise introduced for privacy, maintaining the efficiency and privacy of the centralized version but adapting it for distributed environments.
@@
However, our privacy guarantees rely on the honest-but-curious threat model and assume no collusion or malicious behavior from participants, which may not always be realistic. In the future, it would also be interesting to analyze a private and decentralized version of the accelerated version of the randomized power method, which could potentially have a faster convergence.
8 Impact Statement (required by ICML)
Our contributions directly address the need for privacy-preserving algorithms in the context of large-scale machine learning systems, where sensitive personal or organizational data is often at stake. The decentralized implementation ensures that the method can be applied in some real-world environments where data is distributed across multiple users or devices, making it particularly suitable for recommender systems and other distributed applications. This work is expected to have implications for privacy-conscious machine learning in real-world systems, to perform large-scale data analysis while maintaining robust privacy protections.
@@ Appendix A Proofs of Results
$\begin{split}\Delta_{l}&=\sqrt{\sum_{i}\|\bm{C}_{i:}\bm{X}^{l}\|_{F}^{2}}\\ &\leq\sqrt{\sum_{i}(\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}(\cdot\sum_{j}|\bm{C}_{ij}|)^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\sqrt{\sum_{i}\|\bm{C}_{i:}\|_{1}^{2}}\\ &\leq\max_{i}\|\bm{X}_{i:}^{l}\|_{F}\triangleq\hat{\Delta}_{l}.\\ \end{split}$ (24)
Although the proposed update model (12) is more general than when $\bm{A}^{\prime}$ is defined using (11), the proposed bound $\hat{\Delta}_{l}$ is also generally tighter than the bound proposed in (Hardt & Price, 2014; Balcan et al., 2016).
A.2 Privacy proof
A.2.1 Note on related privacy proofs:
Several Differential Privacy (DP) proofs for the private randomized power method (PPM) have been developed in prior works, for instance those by Hardt & Price (2014), Balcan et al. (2016), Hardt & Roth (2013), Hardt & Roth (2012). These papers mostly rely on the same proofs to establish that the PPM satisfies ($\epsilon$, $\delta$)-DP. Specifically, Balcan et al. (2016) references the privacy proof from Hardt & Price (2014), which builds upon the privacy proof from Hardt & Roth (2013). The proof therein is also closely related to the one of Hardt & Roth (2012).
However, both Hardt & Roth (2013) (Theorem 2.4) and Hardt & Roth (2012) (Theorem 2.4) contain errors in their proposed composition rules, where a comparison sign is mistakenly flipped. This error could potentially cause the privacy parameter $\epsilon$ in the proposed mechanism to be arbitrarily small, providing no privacy guarantee at all. The original composition rule is presented in Theorem III.3 of Dwork et al. (2010). Moreover, Lemma 3.4 of Hardt & Roth (2013) misuses Theorem 2.4. Indeed, they claim that their algorithm satisfies $(\epsilon^{\prime},\delta)$-DP at each iteration. By their Theorem 2.4, then the algorithm overall should satisfy $(\epsilon^{\prime},k\delta+\delta^{\prime})$-DP where $\delta^{\prime}>0$, and not $(\epsilon^{\prime},\delta)$-DP as claimed.
A.2.2 Proof of Theorem 3.2
@@ A.2.1 Note on related privacy proofs:
$\leq\sum_{i=1}^{L}\frac{(\hat{\Delta}_{l})^{2}}{2\sigma_{i}^{2}}$ (26)
@@
By Proposition 1.3 of Bun & Steinke (2016), if M provides $\rho$-zCDP, then M is ($\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP, $\forall\delta>0$. Let $\epsilon^{\prime}\triangleq\rho+2\sqrt{\rho\log(1/\delta)}$.
@@
Which leads us to the following Theorem by injecting $\sigma_{l}=\hat{\Delta}_{l}\cdot\epsilon^{-1}\sqrt{4L\log(1/\delta)}$.
A.4 Derivation of Approximation 6.2
We provide here the derivation for Approximation 6.2:
@@
Let $\bm{M}\in\mathbb{R}^{n\times k}$ with i.i.d $\mathcal{N}(0,1)$-distributed entries. Let $\bm{M}=\bm{Q}\bm{R}$ be its QR factorization (by definition $\bm{i}sQ$ orthogonal and $\bm{R}$ upper triangular). By the Barlett decomposition theorem (Muirhead, 2009), we know that $\bm{Q}$ is a random matrix distributed uniformly in the Stiefel manifold $\mathbb{V}_{k,n}$. Then by Theorem 2.2.1 of Chikuse (2012), we know that a random matrix $\bm{Q}$ uniformly distributed on $\mathbb{V}_{k,n}$ can be expressed as $\bm{Q}=\bm{Z}(\bm{Z}^{\top}\bm{Z})^{-\frac{1}{2}}$ with $\bm{Z}$ another matrix with i.i.d $\mathcal{N}(0,1)$-distributed entries. We approximate $\bm{Z}^{\top}\bm{Z}$ as a diagonal matrix and remark that its diagonal elements are distributed as random chi-squared variables with $n$ degrees of liberty. For a matrix $\bm{A}$, we note as $\bm{A}^{\circ n}$ the elementwise Hadamard exponentiation. Then we have:
@@
We can approximate the expectancy of the maximum (noted as m) of $d$ Gaussian variables distributed as $\mathcal{N}(\mu_{2},\sigma_{2}^{2})$ by $m=\sigma_{2}\sqrt{2\cdot\log(d)}+\mu_{2}$ by Lemma 2.3 from Massart (1896). We therefore get approximations of $\mathbb{E}[\max_{ij}k\bm{Q}_{ij}^{2}]$ and $\mathbb{E}[\max_{i}\|\bm{Q}_{i:}\|^{2}]$. We assume that the variances of $\Delta_{l}^{prior}$ and $\hat{\Delta}_{l}$ (as $n$ and $k$ are large) are small and therefore:
$\mathbb{E}[\Delta_{l}^{prior}]^{2}\approx\mathbb{E}[(\Delta_{l}^{prior})^{2}]=\sqrt{\frac{k^{2}\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(kn)}+\frac{k\mu}{n}$ (46)
$\mathbb{E}[\hat{\Delta}_{l}]^{2}\approx\mathbb{E}[(\hat{\Delta}_{l})^{2}]=\sqrt{\frac{k\sigma^{2}}{n^{2}}}\sqrt{2\cdot\log(n)}+\frac{k\mu}{n}.$ (47)
@@
First bound: We know by Section 2.4 from Woodruff et al. (2014) that any row $\bm{v}$ of a matrix with orthonormal columns $\bm{Z}$ has squared norm at most 1. Therefore $\max_{i}\|\bm{X}_{i:}\|_{2}\leq 1$ because $\bm{X}$ has orthonormal columns by construction.
Second bound: Without loss of generality and to simplify notation, we note $\bm{X}$ to denote a matrix $\bm{X}_{l}$ computed during the PPM. As proposed in Hardt & Price (2014), we can write any column $x_{c}$ of $\bm{X}$ ($\bm{X}_{:c}$) as $\sum_{i=1}^{n}\alpha_{c}^{i}u_{i}$ where $u_{i}$ denotes the $i$-th eigenvector of $\bm{A}$, and $\alpha_{c}^{i}$ is a scalar. Therefore we have: